Shift Summary for owl shift, 06.16.99

Jim Graham (jagraham@hep.uchicago.edu)
Wed, 16 Jun 1999 07:27:29 -0500

Shift Summary for owl shift, 06.16.99

On Shift: JAG, LM

Hours of beam delivered at all intensities: 8.0
Hours of beam delivered at nominal intensity: 8.0
Hours of E832 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.25 hr

Accounting of lost time: pipeline inits.

========================================================================
Runs ended this shift:
Run 13615 - 2.2M events - BEAM_832 - beam studies
Run 13618 - 571K events - BEAM_832 - DC operating point studies
Run 13619 - 620K events - BEAM_832 - " " " "
Run 13621 - 800K events - MUON_4E12-
Run 13623 - still running.

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Beam studies
DC operating voltage studies.

========================================================================
Description of Shift:

run spill time comment
- ------------------------------------------------------------------------
13615 00:37 Inherit run from eve shift. Have pipeline ORD
number and Sparse code errors ~every other spill.
Need to stop triggers, reload pipeline. Seems to
be related to large events in stream 2 (see log p.246)

02:13 Rick Ford ends beam studies. Had beam 0.75mm west
on target. East/West Momentum spectrum still flat.
He returns beam to nominal position.

This run was also used for DC studies with DC 1,2
cathodes = 2400 V, window @ 2300 V.

Kessler remakes L3 for the first time this shift.

13618 1 02:44 Start more DC studies. DC 1,2 2400 V, window 2200 V.
03:22 End run.

13619 03:25 DC studies DC 1,2 2450 V, window 2450 V.

03:35 In an attempt to alleviate the need to reset the
pipeline, we add 50 us dummy busy in ktrig.config.
Livetime drops from 86% to 68%.

18 03:49 Multiple stream 5 kfrend problems fixed by cycling
laser repeater multiple times and kfrending mult.
times.

04:23 End run. Kessler remakes filter code for second
time

SUMMARY of DC STUDIES:
- At cathode = window = 2350 V, isolated singles rate is reduced
by 20%, but with worse efficiency and high-sod rate.
- with V(cat) - V(win) = 100, 200 V, the isolated singles rate is
reduced by at least 10%, but eff. and high-sods still good.
- singles rate depends strongly on beam rate and quality.
- more to follow.

13621 04:25 Start muon run. Elizabeth finds that peds downloaded
from run 13610 are off by roughly 5 counts for the
regenerator.

13623 05:08 Start E832 run.

05:38 Change DUMMY busy back to 8 us from 50, because
Rick K. finds a possible problem with the ETCOS
calibration trigger. Beam events get into this
trigger off-spill, indicating that there is a
problem with the spill signal, either in Ktev,
or wrt beam. However, we don't understand why
this causes pipeline sparse code problems.

BUT, after disabling this trigger, AND changing
the dummy busy, we don't have to reload the pipeline
for the rest of the shift.

========================================================================
Follow up Notes on Previous Shift

- didn't get to the requested laser scan

========================================================================
Items to Follow up on Next Shift

- beam events offspill?
- do laser scan if beam goes away.

Regards, Jim

- --HAA19399.929535907/hep.uchicago.edu--

------- End of Forwarded Message